OpenStudy (anonymous):

The geometric mean of 8 and 2 is 4. True or False?

OpenStudy (anonymous):

true... geometric mean = \(\sqrt{8 \cdot 2} \)

OpenStudy (ajprincess):

Geometric mean of a and b is \(\sqrt{a*b}\)